OpenStudy (anonymous):

definite integral: [(ln(x))^4 / 2x ]dx upper limit e^2 and lower limit 1. Do I use substitution and how do I do this?

ganeshie8 (ganeshie8):

sub u = lnx du = 1/x dx

zepdrix (zepdrix):

Ya a u-sub will work really nicely for this one, maybe this picture will help a little bit. I rearranged things so you can see how your u-sub will get plugged in.

OpenStudy (anonymous):

Wow, once you know what you're doing it becomes VERY easy! Wish I had more time to practice but unfortunately time is limited :( Thanks so much for your help! Once i replaced (1/x dx) with du and integrated 1/2 u^4 the answer is ln^5/10 + C

zepdrix (zepdrix):

Cool, looks good! :)

OpenStudy (anonymous):

When I sub in the limits I get [ln(e^2)^5) - ln(1)^5/10) lne^10 - ln^5/10 does this make sense?

zepdrix (zepdrix):

The exponent ( the 5 ) is being applied to the log, not to the contents of the log. First simply ln(e^2) and ln(1) before you apply the 5th power to them.

OpenStudy (anonymous):

ok that makes sense lne^2=2 and ln(1) is 0 so then it becomes (2)^5/10 - (0)^5/10 = 5/2

zepdrix (zepdrix):

Mmm ya that looks good!

OpenStudy (anonymous):

Appreciate the help!

zepdrix (zepdrix):

Ooo wolfram is coming up with 32/5.. Lemme see where we goofed up..

zepdrix (zepdrix):

Err 16/5* hmm

OpenStudy (anonymous):

Yep I got it.

zepdrix (zepdrix):

Hmm your answer looks correct, I don't know where 16/5 is coming from.. weirdddd :U

OpenStudy (anonymous):

32/10 = 16/5 I think it's ok. My answer must have been incorrect as (2)^5 = 32